# The Unbounded Denominators Conjecture

### v1

@@ 2.4. The extrapolation
This completes the proof of Theorem 2.0.2. ∎
@@
André pointed out to us that the alternative version
(2.4.3) $m\leq\frac{\Big(1+\frac{1}{\kappa}\Big)\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}+\kappa C}{\log{|\varphi^{\prime}(0)|}}$
of (2.1.1), where $C\in\mathbf{R}$ is the constant in Lemma 2.1.2 (2) and the parameter $\kappa>0$ is arbitrary, can be more directly proved using a single variable auxiliary construction and the subharmonic property of $\log{|z^{-\beta}F(\varphi(z))|}$, instead of our use of the maximum modulus principle in a high number $d\to\infty$ of variables. In the specificity of our current application, it turns out that we will have $C=O(\log{N})$, $\int_{\mathbf{T}}\log^{+}{|p\circ\varphi|}\,\mu_{\mathrm{Haar}}=O(\log{N})$, and $\log{|\varphi^{\prime}(0)|}\gg N^{-3}$. Thus André’s variant (2.4.3) with the choice $\kappa=1$ is also sufficient for the proof of the unbounded denominators conjecture.
Remark 2.4.4.
Another related approach, also suggested to us by André, is to use an induction on the number of variables and the plurisubharmonic property of $\log{|z_{1}^{-n_{1}}F(\varphi(\mathbf{z}))|}$, where $n_{1}$ is the lowest partial degree of $x_{1}$ in $F(\mathbf{x})$, instead of the maximum modulus principle for $V(\mathbf{z})^{M}F(\varphi(\mathbf{z}))$. Such a procedure turns out to work well if we consider, in place of the leading order jet, instead the overall (possibly of a higher total degree $|\mathbf{n}|>\beta$) lexicographically lowest monomial $c\,\mathbf{x^{n}}$ in $F(\mathbf{x})$. Indeed this path leads to an alternative, and arguably simpler proof of Theorem 2.0.2. We plan to give further details in a subsequent work and explore the added flexibility of the lexicographically lowest monomial. (We have kept the original argument with the thought that it may have ideas useful in other settings.)
Remark 2.4.5.
The result is more general, and the restriction here to $\mathbf{Z}\llbracket t\rrbracket$ expansions was chosen as minimal for our application to noncongruence modular forms. In a sequel work we will generalize our integrated holonomy rank bound, in particular to the case of $\mathbf{Q}\llbracket t\rrbracket$ formal functions, and study its applications to transcendence theory. With regard to the latter, it is of some interest to inquire about the optimal numerical constant that could take the place of the coefficient $e$ in (2.0.4).
@@ 3. Some reductions
Then the ring $R_{2N}$ of modular functions with bounded denominators and cusp widths dividing $2N$ has dimension at most $CN^{3}\log{N}$ over the ring of modular functions of level $\Gamma(2)$, for some absolute constant $C$.
@@
with the Kummer integrality condition $p(x)=x^{N}\in\mathbf{Z}\llbracket q\rrbracket=\mathbf{Z}\llbracket t^{N}\rrbracket\subset\mathbf{Z}\llbracket t\rrbracket$ being in place. The integrality and cusp widths conditions in the definition of the ring $R_{2N}$ entail that for all $f\in R_{2N}$, the formal expansions $f\in\mathbf{Q}\llbracket x\rrbracket$ belong to the ring $\mathcal{H}(U,x(t),\mathbf{Z})\otimes_{\mathbf{Z}}\mathbf{Q}$. It thus suffices to bound the $\mathbf{Q}[x^{N}]$-dimension of the latter ring by $CN^{3}\log{N}$.
We take $r:=1-AN^{-3}/2$ and
@@
Theorem 2.0.2 now upper-bounds our requisite dimension as
$\dim_{\mathbf{Q}(x^{N})}\big(\mathcal{H}(U,x(t),\mathbf{Z})\otimes\mathbf{Q}(x^{N})\big)\leq e\cdot\frac{\int_{|z|=1-A/(2N^{3})}\log^{+}{|F_{N}^{N}|}\,\mu_{\mathrm{Haar}}}{cN^{-3}}.$
The bound by $O(N^{3}\log{N})$ now results follows from (2) with the choice $B:=A/2$. ∎
We prove both of the required statements in Theorems 5.3.8 and Theorem 6.0.1 respectively. We then put up the $CN^{3}\log{N}$ dimension bound supplied by Proposition 3.0.1 for the ring $R_{N}$ of all modular functions against the obvious $\gg N^{3}$ lower bound for the subring of the congruence examples from the fact that $\Gamma(2):\Gamma(2N)]\gg N^{3}$ (see Equation 4.2.2). Hence we have:
@@ 3.1. A summary
    In §5, we study the properties of the function $F_{N}$. It turns out more or less to be related to a Schwarzian automorphic function on a (generally noncongruence) triangle group. This allows us to compute the conformal radius of $F_{N}$ exactly (see Theorem 5.3.8), and indeed it has the form $16^{1/N}\big(1+(\zeta(3)/2)N^{-3}+\cdots\big)$.
@@ 4.1. Wohlfahrt Level
We being by recalling a notion of level for noncongruence subgroups due to Wohlfahrt [Woh64]. Let $G\subset\mathrm{SL}_{2}(\mathbf{Z})$. If $\gamma\in G$ is any parabolic element, then $\gamma$ is conjugate in $\mathrm{SL}_{2}(\mathbf{Z})$ to a power $U^{m}$ of the matrix
@@
The level $L(G)$ of $G$ is the lowest common multiple of $m$ as $\gamma$ ranges over all parabolic $\gamma$ which generate the stabilizer of some cusp of $G$ — equivalently, over all parabolic $\gamma$ which are not non-trivial powers of some other parabolic element.
We begin with some elementary properties concerning this definition. We typically only consider groups containing I-I since we are secretly interested in subgroups of $\mathrm{PSL}_{2}(\mathbf{Z})$.
@@
If $G$ and $H$ both contain I-I and have $L(G)$ and $L(H)$ dividing $N$, then any cusp of $G\cap H$ also has cusp width dividing $N$.
@@
The stabilizer of a cusp inside any subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$ containing I-I is $\mathbf{Z}/2\mathbf{Z}\times\mathbf{Z}$. In particular, if $G$ contains the group $a\mathbf{Z}$ and $H$ contains $b\mathbf{Z}$ then $G\cap H$ contains $\mathrm{lcm}(a,b)\mathbf{Z}$, and the result follows. ∎
Corollary 4.1.4.
Let $G\subset\mathrm{SL}_{2}(\mathbf{Z})$ be a finite index subgroup containing I-I with Wohlfahrt level $N$. Let $\widetilde{G}$ be the intersection of the conjugates of $G$ by $\mathrm{SL}_{2}(\mathbf{Z})$, so $\widetilde{G}$ is a normal subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$ contained in $G$. Then $\widetilde{G}$ has level $N$.
@@
Let $H\subset\mathrm{SL}_{2}(\mathbf{Z})$ have $L(H)=N$. Then $\widetilde{H}:=A^{-1}HA\cap\mathrm{SL}_{2}(\mathbf{Z})$ has L(H~)L($\widetilde${H}) dividing $Np$.
@@
Any unipotent element in $\widetilde{H}$ is conjugate by $A$ to a unipotent element in $H$, which is a power of a minimal unipotent element $\gamma\in H$. We may write
@@
with $n|N$, and ${B=\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\in\mathrm{SL}_{2}(\mathbf{Z})}$. We have
$\widetilde{\gamma}:=(A^{-1}BA)\left(\begin{matrix}1&n/p\\ 0&1\end{matrix}\right)(A^{-1}BA)^{-1},$
where ${(A^{-1}BA)=\left(\begin{matrix}a&b/p\\ cp&d\end{matrix}\right)}$. We consider two cases:
@@
    Suppose that $(a,p)=1$. Since $(a,c)=1$ we have $(a,pc)=1$. Hence we may write
@@
    where
    $C=\left(\begin{matrix}a&e\\ cp&f\end{matrix}\right)\in\mathrm{SL}_{2}(\mathbf{Z})\text{ such that }A^{-1}BA=C\left(\begin{matrix}1&h/p\\ 0&1\end{matrix}\right),\,h\in\mathbf{Z},$
    for any suitable choice of $e,f$ and $h$. We now have
@@
    and thus in $\widetilde{H}$, and hence the cusp width at this cusp divides $N$ and hence also $Np$.
@@
    and so the cusp width divides $Np$. ∎

### v2

@@ 2.5.3. The extrapolation
This concludes also our original proof of Theorem 2.0.1. ∎
2.5.4. A potential-theoretic generalization
The path with §§2.1,2.5 leads straightforwardly to an extension in potential theory, which we formulate without detailing a proof. Consider a compact subset $K\subset\overline{D(0,1)}$ with transfinite diameter $d(K):=\mathrm{cap}(K,[\infty])$. The equilibrium measure $\mu_{K}$ is the unique probability measure supported by $K$ that attains the lowest possible logarithmic energy of
$\iint_{K\times K}\log{\frac{1}{|x-y|}}\,\mu_{K}(x)\,\mu_{K}(y)=-\log{d(K)}.$
If
$\log{|\varphi^{\prime}(0)|}+\log{d(K)}>0,$
then under the hypotheses of Corollary 2.0.5 we have the holonomy rank bound
(2.5.24) $\dim_{\mathbf{Q}(p(x))}\mathcal{V}(U,x(t),\mathbf{Z})\leq e\,\frac{\int_{K}\log^{+}{|p\circ\varphi|}\,\mu_{K}}{\log{|\varphi^{\prime}(0)|}+\log{d(K)}}.$
The cases $K=\overline{D(0,1)}$ or $K=\mathbf{T}$ both recover Corollary 2.0.5.
@@ 2.5.4. A potential-theoretic generalization
The result is still more general than 2.5.4, and the restriction here to $\mathbf{Z}\llbracket t\rrbracket$ expansions was chosen as minimal for our application to noncongruence modular forms. In a sequel work we will generalize our integrated holonomy rank bound, in particular to the case of $\mathbf{Q}\llbracket t\rrbracket$ formal functions, and study its applications to transcendence theory. With regard to the latter, it is of some interest to inquire about the optimal numerical constant that could take the place of the coefficient $e$ in (2.5.24).
@@ 3. Some reductions
Then the $\mathbf{Q}(\lambda)$-vector space $R_{2N}$ generated by the modular functions with Fourier coefficients in $\mathbf{Q}$ and bounded denominators at the cusp $\zeta=i\infty$, and having cusp widths dividing $2N$ at all cusps $\zeta\in\mathbf{P}^{1}(\mathbf{Q})$, has dimension at most $CN^{3}\log{N}$ over the field $\mathbf{Q}(\lambda)$ of modular functions of level $\Gamma(2)$, for some absolute constant $C$.
@@
with the Kummer integrality condition $p(x)=x^{N}\in\mathbf{Z}\llbracket q\rrbracket=\mathbf{Z}\llbracket t^{N}\rrbracket\subset\mathbf{Z}\llbracket t\rrbracket$ being in place.
The integrality and cusp widths conditions in the definition of the $\mathbf{Q}(\lambda)$-vector space $R_{2N}$ entail a basis of $R_{2N}$ made of elements of the ring $\mathcal{H}(U,x(t),\mathbf{Z})\otimes_{\mathbf{Z}}\mathbf{Q}$. More precisely, for a modular function $f$ with Fourier expansion at $i\infty$ lying in $\mathbf{Z}\llbracket q^{1/N}\rrbracket\otimes_{\mathbf{Z}}\mathbf{Q}$, by our choice of $t=q^{1/N}$ and $x(t)=(\lambda(q)/16)^{1/N}$ we have $x^{*}f\in\mathbf{Z}\llbracket t\rrbracket\otimes_{\mathbf{Z}}\mathbf{Q}$, on defining $x^{*}f$ as the formal $x$-expansion of $f(q)=f(q(x))\in\mathbf{Z}\llbracket q^{1/N}\rrbracket\otimes_{\mathbf{Z}}\mathbf{Q}\subset\mathbf{Q}\llbracket x\rrbracket$. As $f$ is a regular function on some affine modular curve $Y$ over $\overline{\mathbf{Q}}$ which admits a finite étale map to $Y(2)_{\overline{\mathbf{Q}}}$, there exists a (minimal) nonzero linear differential operator $L$ over $\overline{\mathbf{Q}}(\lambda)$ such that $L(f)=0$ and all singularities $\lambda\neq 0,1,\infty$ of $L$ in $\mathbf{C}\smallsetminus\{0,1\}$ have trivial local monodromy. Moreover, our assumption on the cusp widths dividing $2N$ implies that a local coordinate in a small neighborhood of each cusp of $Y$ above $\lambda=0$ can be chosen to be the lift of some (positive integer) power of $x=(\lambda/16)^{1/N}$. This means that the pullback of $L$ to $U\smallsetminus\{0\}=\mathbf{C}^{\times}\smallsetminus 16^{-1/N}\mu_{N}$ admits a full set of meromorphic solutions in some sufficiently small neighborhood of $x=0$, i.e. has a trivial local monodromy around $x=0$. Therefore $f\in\mathcal{H}(U,x(t),\mathbf{Z})\otimes_{\mathbf{Z}}\mathbf{Q}$, and $R_{2N}\subset\mathcal{V}(U,x(t),\mathbf{Z})$.
It thus suffices to bound $\dim_{\mathbf{Q}(x^{N})}\mathcal{V}(U,x(t),\mathbf{Z})$ by $CN^{3}\log{N}$. We take $r:=1-AN^{-3}/2$ and
@@
Corollary 2.0.5 now upper-bounds our requisite dimension as
$\dim_{\mathbf{Q}(x^{N})}\mathcal{V}(U,x(t),\mathbf{Z})\leq e\cdot\frac{\int_{|z|=1-A/(2N^{3})}\log^{+}{|F_{N}^{N}|}\,\mu_{\mathrm{Haar}}}{cN^{-3}}.$
The requisite bound by $O(N^{3}\log{N})$ now results follows from (2) with the choice $B:=A/2$. ∎
Remark 3.0.5.
We may also prove this proposition by using Theorem 2.0.1 directly. Using the notation in the proof, let $Y^{\prime}$ denote the modular curve $Y$ with all the cusps above $0\in Y(2)$ filled in. The fiber product $Y^{\prime}\times_{Y(2)_{\overline{\mathbf{Q}}}}U$ with its natural map to $U$ is a covering map (one can check this claim locally; the assumption on cusp widths is used to prove that $0\in U$ is not ramified). Therefore, the universal covering map $D(0,1)\rightarrow U$ factors through $Y^{\prime}\times_{Y(2)_{\overline{\mathbf{Q}}}}U$ and thus we obtain a map $D(0,1)\rightarrow Y^{\prime}\times_{Y(2)_{\overline{\mathbf{Q}}}}U\rightarrow Y^{\prime}$ (the second map is the natural map) such that its composition with $Y^{\prime}\rightarrow Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}$ is the map $D(0,1)\rightarrow U\rightarrow Y(2)_{\overline{\mathbf{Q}}}\cup\{0\}$. Thus $f\circ\varphi$ is also given by the natural pullback of $f$ from $Y^{\prime}$ to $D(0,1)$ and thus it is holomorphic over $D(0,1)$ as far as $f$ is holomorphic at all cusps in $Y^{\prime}$, which can be achieved by multiplying $f$ with a suitable power of $\lambda$. Thus we verify the analyticity property in Theorem 2.0.1. The rest of the proof is the same as the proof above.
We prove both of the required statements in Theorems 5.1.4 and Theorem 6.0.1 respectively. We then put up the $CN^{3}\log{N}$ dimension bound supplied by Proposition 3.0.1 for the vector space $R_{2N}$ of all modular functions against the obvious $\gg N^{3}$ lower bound for the subring of the congruence examples from the fact that $\Gamma(2):\Gamma(2N)]\gg N^{3}$ (see Equation 4.2.4). Hence we have:
@@ 3.1. A summary
    In §5, we study the properties of the function $F_{N}$. It turns out more or less to be related to a Schwarzian automorphic function on a (generally non-arithmetic) triangle group. This allows us to compute the conformal radius of $F_{N}$ exactly (see Theorem 5.1.4), and indeed it has the form $16^{1/N}\big(1+(\zeta(3)/2)N^{-3}+\cdots\big)$.
@@ 4.1. Wohlfahrt Level
We begin by recalling a notion of level for noncongruence subgroups due to Wohlfahrt [Woh64]. Let $G\subset\mathrm{SL}_{2}(\mathbf{Z})$ be a finite index subgroup. (Many of the arguments of this section do not require this hypotheses but since it is satisfied for our applications we assume it to avoid unnecessary distractions.) The group consisting of the two matrices $\pm I$, where $I=\left(\begin{matrix}1&0\\ 0&1\end{matrix}\right)$, will be denoted by $E$. The group $\mathrm{SL}_{2}(\mathbf{Z})$ acts via Möbius transformations both on the upper half plane $\mathbf{H}$ and the extended upper half plane $\mathbf{H}^{*}=\mathbf{H}\cup\mathbf{P}^{1}(\mathbf{Q})$. The action of $\mathrm{SL}_{2}(\mathbf{Z})$ on $\mathbf{P}^{1}(\mathbf{Q})$ is transitive. It follows that if a non-trivial element $\gamma\in G$ fixes an element $\zeta\in\mathbf{P}^{1}(\mathbf{Q})$, then $\gamma$ has the form $\pm MU^{m}M^{-1}$, where $M\in\mathrm{SL}_{2}(\mathbf{Z})$, $M\infty=\zeta$, and
@@
We call such a $\zeta\in\mathbf{P}^{1}(\mathbf{Q})$ a cusp of $G$. If $\zeta\in\mathbf{P}^{1}(\mathbf{Q})$, then $MU^{m}M^{-1}=(MUM^{-1})^{m}\in G$ for some $m$ because $G$ has finite index in $\mathrm{SL}_{2}(\mathbf{Z})$, and hence every element of $\mathbf{P}^{1}(\mathbf{Q})$ is a cusp of $G$. The stabilizer in $G$ of a cusp $\zeta\in\mathbf{P}^{1}(\mathbf{Q})$ is either isomorphic to $E\times\mathbf{Z}=\mathbf{Z}/2\mathbf{Z}\times\mathbf{Z}$ or $\mathbf{Z}$, depending on whether $E\subset G$ or not. For each $\zeta$, there is a minimal positive integer $m$ such that $\pm MU^{m}M^{-1}\in G$, and we say that $m$ is the width of the cusp $\zeta$. The action of $G$ on $\mathbf{P}^{1}(\mathbf{Q})$ has finitely many orbits, and the cusp width only depends on the orbit of the cusp under $G$. Geometrically, the complex structure on $\mathbf{H}$ imbues the quotient $X(G)=\mathbf{H}^{*}/G$ with the structure of an algebraic curve. From this point of view, the equivalence classes of cusps of $G$ (up to the action of $G$) are in bijection with the pre-images of $\infty$ under the projection $X(G)\rightarrow X(\mathrm{SL}_{2}(\mathbf{Z}))=\mathbf{P}^{1}_{j}$, and the cusp widths are exactly the ramification indices of this map at $j=\infty$.
@@
The level $L(G)$ of $G$ is the lowest common multiple of all the cusp widths of $G$.
We begin with some elementary properties concerning this definition. We typically only consider groups containing $E=\langle-I\rangle$ since we are generally interested in stabilizers of functions under Möbius transformations.
@@
Let $G$ and $H$ be finite index subgroups of $\mathrm{SL}_{2}(\mathbf{Z})$ both containing $E$. Suppose that $L(G)$ and $L(H)$ both divide $N$, then any cusp of $G\cap H$ also has cusp width dividing $N$.
@@
The stabilizer of a cusp inside any subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$ containing $E$ is $E\times\mathbf{Z}$. In particular, if $G$ contains the group $E\times a\mathbf{Z}$ and $H$ contains $E\times b\mathbf{Z}$ then $G\cap H$ contains $E\times\mathrm{lcm}(a,b)\mathbf{Z}$, and the result follows. ∎
Lemma 4.1.4.
Let $G\subset\mathrm{SL}_{2}(\mathbf{Z})$ be a finite index subgroup containing $E$ with Wohlfahrt level $N$. Let $N(G)$ be the largest normal subgroup of $\mathrm{SL}_{2}(\mathbf{Z})$ contained in $G$. Then $N(G)$ has finite index in $\mathrm{SL}_{2}(\mathbf{Z})$ and $L(N(G))=N$.
Proof.
Since $G$ has finite index in $\mathrm{SL}_{2}(\mathbf{Z})$, the group $N(G)$ is the intersection of the finitely many conjugates of $G$ by $\mathrm{SL}_{2}(\mathbf{Z})$. Hence $N(G)$ has finite index and $L(N(G))=N$ by Lemma 4.1.3. ∎
@@
Let $H\subset\mathrm{SL}_{2}(\mathbf{Z})$ be a finite index subgroup containing $E$ such that $L(H)=N$. Then $L(A^{-1}HA\cap\mathrm{SL}_{2}(\mathbf{Z}))$ divides $Np$.
@@
Let us write $\widetilde{H}:=A^{-1}HA\cap\mathrm{SL}_{2}(\mathbf{Z})$. Note that $\widetilde{H}$ contains $A^{-1}EA=E$. In particular, the stabilizer of any cusp of $\widetilde{H}$ has the form $E\times\mathbf{Z}$, where the $\mathbf{Z}$ is generated by a unipotent element $\widetilde{h}$ of $\widetilde{H}$ conjugate in $\mathrm{SL}_{2}(\mathbf{Z})$ to $U^{m}$ for some positive integer $m$, and we want to show that $m$ divides $Np$.
Any unipotent element $\widetilde{h}$ in $\widetilde{H}$ has the form $\widetilde{h}=A^{-1}hA$ for some unipotent element $h\in H$. The element $h\in H$ will stabilize some cusp $\zeta\in\mathbf{P}^{1}(\mathbf{Q})$. The stabilizer of $\zeta$ in $H$ has the form $E\times\mathbf{Z}$ where $\mathbf{Z}$ is generated by a unipotent element $\gamma\in H$. It follows that $h$ will be the smallest power of $\gamma$ which lies in $\widetilde{H}$, or equivalently in $\mathrm{SL}_{2}(\mathbf{Z})$. Since $L(H)=N$, we may write
@@
with $n|N$, and ${B=\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\in\mathrm{SL}_{2}(\mathbf{Z})}$. We define
$\widetilde{\gamma}=A^{-1}\gamma A=$ $A^{-1}B\left(\begin{matrix}1&n\\ 0&1\end{matrix}\right)B^{-1}A$
$=$ $\ (A^{-1}BA)\left(A^{-1}\left(\begin{matrix}1&n\\ 0&1\end{matrix}\right)A\right)(A^{-1}B^{-1}A)$
$=$ $\ (A^{-1}BA)\left(\begin{matrix}1&n/p\\ 0&1\end{matrix}\right)(A^{-1}BA)^{-1},$
where ${(A^{-1}BA)=\left(\begin{matrix}a&b/p\\ cp&d\end{matrix}\right)}$. Since $h$ is a power of $\gamma$, we deduce that $\widetilde{h}=A^{-1}hA$ is a power of $\widetilde{\gamma}$, although $\widetilde{\gamma}$ need not be in $\widetilde{H}$ since it is not necessarily integral. We consider two cases:
@@
    Suppose that $(a,p)=1$. Since $(a,c)=1$ we have $(a,pc)=1$, and thus there exist $r,s\in\mathbf{Z}$ with asrpc=1as-rpc=1, and hence
    $C=\left(\begin{matrix}a&r\\ pc&s\end{matrix}\right)\in\mathrm{SL}_{2}(\mathbf{Z}).$
    For such a $C$, we have, with t=bsdprt=bs-dpr, the identity
    $A^{-1}BA=C\left(\begin{matrix}1&t/p\\ 0&1\end{matrix}\right).$
    But since $\left(\begin{matrix}1&t/p\\ 0&1\end{matrix}\right)$ commutes with $\left(\begin{matrix}1&n/p\\ 0&1\end{matrix}\right)$, it follows that we may write
@@
    We now have
@@
    and thus in $\widetilde{H}$. Hence either $\widetilde{h}=\widetilde{\gamma}$ if $\widetilde{\gamma}$ lies in $\mathrm{SL}_{2}(\mathbf{Z})$ or $\widetilde{h}=(\widetilde{\gamma})^{p}$. In particular, the cusp width at this cusp is either $n$ or $n/p$ and certainly divides $N$ and hence also $Np$.
@@
    and hence $\widetilde{h}=\widetilde{\gamma}$ and the cusp width at this cusp is $np$ which divides $Np$. ∎